OpenStudy (unklerhaukus):

Evaluate \[\int\limits_{0}^{1}[{{1-x^2 \over 1+x^2}]^{1/2}}dx\] In terms of the Gamma function

OpenStudy (anonymous):

hello myininaya!

OpenStudy (anonymous):

hope you had a good holiday.

OpenStudy (unklerhaukus):

hmm

OpenStudy (anonymous):

does it involve trig substitution

OpenStudy (anonymous):

could you please tell me in which chapter this in calc

OpenStudy (unklerhaukus):

when i tried trig substitution 1-x^2 = cos^2(x) 1+x^2 = sec^2(x) i get \[\int\limits_{0}^{1}\cos^2(x) dx\]

OpenStudy (anonymous):

use the half angle formula

OpenStudy (unklerhaukus):

subject; differential equations In my book the chapter is; Special Functions; the topic is; beta and gamma functions

OpenStudy (unklerhaukus):

i can not see where to apply the half angle formula salem123

OpenStudy (anonymous):

cos^2x=1/2(1+cos2x)

OpenStudy (anonymous):

do u think this works?

OpenStudy (unklerhaukus):

i dont know

OpenStudy (anonymous):

but how did you end up only with cos(x)^2

OpenStudy (anonymous):

how did you get rid off sec(x)^2

OpenStudy (unklerhaukus):

well √cos^2(x)/sec^2(x) = √ cos^4(x) = cos^2(x)

OpenStudy (unklerhaukus):

right ?

OpenStudy (anonymous):

okay i see

OpenStudy (anonymous):

right

OpenStudy (anonymous):

if this is right then the half angle formula should be applicable then

OpenStudy (anonymous):

but the i am looking how to express it in term of gamma

OpenStudy (unklerhaukus):

by the way i have the answer \[={√π \over 4} \times ({\Gamma(1/4) \over \Gamma(3/4)}-4{\Gamma(3/4) \over \Gamma(1/4)})\]

OpenStudy (anonymous):

you want to know how to get there

OpenStudy (unklerhaukus):

yeapp

OpenStudy (anonymous):

no idea to be honest i can't help

OpenStudy (anonymous):

i am taking differential equation but we haven't taken this stuff yet

OpenStudy (unklerhaukus):

thanks any way

OpenStudy (jamesj):

This term \( \sqrt{\pi} /4 \) is a multiple of \( \Gamma(1/2) \). This is an interesting question; I'll think about it.

OpenStudy (jamesj):

...tomorrow though. Time for bed to digest all the pie and turkey!

OpenStudy (unklerhaukus):

i gotta go to , yeah i noticed the Gamma of (1/2) but couldn't reverse engineer the problem correctly

OpenStudy (unklerhaukus):

working backwards from \[=√π/4×({Γ(1/4) \over Γ(3/4)}−4{Γ(3/4) \ \over Γ(1/4)})\] \[=1/4[{ Γ(1/2)(Γ(1/4) \over Γ(3/4)}−4{Γ(1/2)Γ(3/4) \over Γ(1/4)}]\] \[= {B_{(1/2,1/4)}\over2} + B_{(-1/2, 3/4)} \]

OpenStudy (jamesj):

Never mind. I see now, \( -2\Gamma(1/2) = \Gamma(-1/2) \)

OpenStudy (unklerhaukus):

\[Γ(z) = Γ(z+1)/z\] \[Γ(1/2) =-1/2 Γ(-1/2)\]

OpenStudy (unklerhaukus):

yeah

OpenStudy (jamesj):

Ok. So this is now doable: show the integral is equal to the sum of those two beta functions.

OpenStudy (jamesj):

btw, isn't \( B(x,y) = \frac{ \Gamma(x) \Gamma(y)}{\Gamma(x+y)} \), so have you got the constants quite right?

OpenStudy (unklerhaukus):

i think i have it right (i havent included all the steps, but wolfram can confirmed equivalence)

OpenStudy (jamesj):

screw Wolfram; it's not infallible. You need to know for yourself that this is right. \[ \Gamma(1/2) = \sqrt{\pi} \] so I think the constants aren't quite right.

OpenStudy (unklerhaukus):

oh wait i mean \[=1/4 \times B(1/2,1/4)+1/2 \times B(−1/2,3/4)\]

OpenStudy (unklerhaukus):

(off by a factor of 2)

OpenStudy (jamesj):

Right.

OpenStudy (unklerhaukus):

so \[\int\limits_{0}^{1}\sqrt {(1-x^2) \over (1+x^2)})dx={1 \over 2}({B_{(1/2,1/4)} \over 2} + B_{(-1/2,3/4)})\]

OpenStudy (unklerhaukus):

\[=\int\limits_{0}^{1}(1-x^2)^{1/2}(1+x^2)^{-1/2}dx\]